OpenStudy (anonymous):

complete the square -3x^2-12x-11

OpenStudy (chaise):

-3x^2-12x-11 -3(x^2+4x+4)+1 Do you need help on how this answer was obtained?

jimthompson5910 (jim_thompson5910):

\[\large -3x^2-12x-11\] \[\large -3\left(x^2+4x+\frac{11}{3}\right)\] \[\large -3\left(x^2+4x+4-4+\frac{11}{3}\right)\] \[\large -3\left((x+2)^2-4+\frac{11}{3}\right)\] \[\large -3\left((x+2)^2+\frac{1}{3}\right)\] \[\large -3(x+2)^2-3\left(\frac{1}{3}\right)\] \[\large -3(x+2)^2-1\] So \[\large -3x^2-12x-11=-3(x+2)^2-1\]

OpenStudy (anonymous):

thanks

OpenStudy (chaise):

Yeah - factorise the middle bit you get: -3(x+2)^2 -1 My bad.